### v1

Abstract
Community-based graph clustering is one of the most popular topics in the analysis of complex social networks. This type of clustering involves grouping vertices that are considered to share more connections, whereas vertices in different groups share fewer connections. A successful clustering result forms densely connected induced subgraphs. This paper studies a specific form of graph clustering problems that can be formulated as semi-assignment problems, where the objective function exhibits block properties. We reformulate these problems as sparse-constrained optimization problems and relax them to continuous optimization models. We apply a quadratic penalty method to the relaxation problem and solve the nonlinear quadratic penalty subproblem with simple box constraints using a projected gradient method based on the active set. Extensive numerical results indicate that our method provides more accurate clustering results for solving graph clustering problems at a faster speed, both for synthetic graphs and real-world network datasets, particularly in large-scale cases.
Keywords:
Graph clustering Network community detection Sparse optimization Quadratic penalty method Projected gradient method Semi-assignment problems
MSC:
90C10 90C20 90C27 05C90
journal: Optimization and Engineering
1 Introduction
The community-based graph clustering problem no1 , also known as network community detection, is one of the most popular topics in the analysis of complex networks. It has widespread applications in various fields, including sociology no2 ; no3 , biology no4 ; no5 ; no6 , and computer science no7 ; no8 , with specific uses in areas such as social media no9 , healthcare no10 , the web no11 ; no12 , and path searches no13 ; no14 . A cluster, also referred to as a community, is a set of vertices that are more densely connected to each other than to the rest of the network. The community-based graph clustering problem involves partitioning the vertex set $V$ of a graph into $k$ non-empty subsets, with many edges within each cluster and relatively few edges between clusters. Successful clustering produces clusters with dense internal connections and sparse links between different clusters no15 ; GC-BM . A visual representation of a network with this type of cluster structure is shown in Fig. 1. It is important to note that this paper focuses on clustering undirected unweighted graphs without self-loops and multiple edges, and we assume that vertices are assigned to non-overlapping clusters.
Figure 1: A schematic representation of a network with cluster structure. In this network, there are three clusters of densely connected vertices (indicated by solid circles), which are represented by yellow, purple, and green dashed circles. The density of connections between these clusters is much lower
Unlike the common Euclidean space, graphs do not exist in the form of coordinates, and the distance between two nodes cannot be directly measured. Consequently, traditional clustering algorithms based on Euclidean space, such as K-means, are not applicable to graph clustering problems no16 ; GC-BM . The methods for addressing graph clustering problems can be mainly categorized into five types: (1) spectral algorithms, which utilize the spectral properties of graphs to detect clusters no17 ; no18 ; (2) statistical inference-based sta-model methods, such as Markov no19 and Bayesian approaches no20 ; (3) dynamics-based methods, which involve running dynamic processes on the network to identify clusters, such as diffusion no21 , spin dynamics no22-1 ; no22 and synchronization no23 ; (4) divisive clustering algorithms, a class of top-down hierarchical methods that recursively partition the graph into clusters s-1 ; no24 ; no25 ; (5) optimization-based methods, where the clustering result is obtained by finding the extreme value of a function in the possible clustering space, such as maximizing modularity no24 (optimizing the modularity quality function) or maximizing similarity no27 . For a comprehensive and detailed overview of methods for solving graph clustering problems, readers are referred to s-1 ; no15 ; s-3 .
Optimization-based methods are characterized by low computational cost and high accuracy. The maximization of the modularity function no24 is one of the most popular approaches for community detection, with the Louvain method designed by Blondel et al. no28 being the most well-known and effective modularity maximization technique. Although these methods offer advantages such as short computation time and the ability to operate without requiring the number of clusters as input parameter GC-BM , they also have drawbacks, including the limitations of observed clusters no29 , the resolution limit problem no30 ; no32 and the degeneracy problem no30 ; cq-1 . Consequently, many researchers have explored alternatives that do not rely on modularity maximization to address graph clustering problems, such as those found in no33 ; no34 ; no35 . Miasnikof et al. GC-BM were the first to extend the distance minimization of the binary quadratic formulas of Fan and Pardalos no27 to the general case of graph clustering problems. They employed the Jaccard distance to reflect connectivity and used a Boltzmann machine heuristic to solve the resulting model, which is presented below
$\min\limits_{x}$ $\sum\limits_{i}\sum\limits_{j>i}\sum\limits_{k}x_{ik}x_{jk}d_{ij}$ (1)
$s.t.$ $\sum\limits_{k}x_{ik}=1,\ \forall i,$
$x_{ik}\in\{0,1\},\ \forall i,k.$
Here $x_{ik}$ is a binary variable that takes the value of $1$ if the vertex $i$ is assigned to the cluster $k$, and $d_{ij}$ represents the Jaccard distance between $i$ and $j$. Note that (1) is essentially a binary programming problem with semi-assignment constraints. Inspired by the approach in no37 to solve hypergraph matching problems, in this paper, we consider a general class of graph clustering problems (see (2)), and propose an efficient approach to solve it.
The contributions of this paper are as follows: (1) We first summarize several graph clustering models that can be uniformly formulated as semi-assignment problems with objective functions that have block properties. (2) For this class of problem, we equivalently reformulate them as sparse-constrained optimization problems. (3) We apply a quadratic penalty method to the relaxation problem, and for the nonlinear quadratic penalty subproblem with simple box constraints, we employ a projected gradient method based on the active set. The special structure of this class of problems allows for the acceleration of the computation in this method. (4) Extensive numerical results indicate that, for synthetic or real-world graphs, this algorithm identifies more accurate clusters at a faster speed, especially in large-scale problems.
The remainder of this paper is organized as follows. In Sect. 2, we propose the semi-assignment optimization model for graph clustering and discuss its properties. In Sect. 3, we investigate the continuous relaxation problem and apply the fast quadratical penalty method to solve it. In Sect. 4, we present extensive numerical results to verify the efficacy of the proposed method. Final conclusions are given in Sect. 5.
2 Semi-assignment optimization model for graph clustering
In this section, we will establish the mathematical model for graph clustering, which can be cased as a semi-assignment optimization problem.
2.1 Optimization model based on semi-assignment constraints
To that end, we consider a graph $G=(V(G),E(G))$, where $V(G)=\{1,2,\cdots,n\}$ is a nonempty set of vertices and $E(G)=\{(i,j)\ |\ i,j=1,\cdots,n,\ i\ {\rm and}\ j\ {\rm is}\ {\rm connected}\}$ is a set of edges. Let the set of all clusters be =$\mathbb{C}=\{C_{1},C_{2},\cdots,C_{K}\}$, where $K$ is the number of clusters. We denote the total number of edges (or vertices) by $|E(G)|$ (or $|V(G)|=n$), and the number of vertices in the cluster $i$ by $n^{(i)}$.
Roughly speaking, graph clustering involves dividing the vertices of a graph into several subsets of densely connected vertices, where vertices within the same subset have more common connections than those in different subsets. Mathematically speaking, let $x_{ik}$ denote whether vertex $i$ is assigned to cluster $k$, and $x_{ik}$ is one if yes and zero otherwise. The essential aim of graph clustering is to assign each vertex a cluster such that some measures of the clustering is minimized. That is,
$\min\limits_{x}$ $f(x)$ (2)
$s.t.$ $\sum\limits_{k=1}^{K}x_{ik}=1,\ i=1,\cdots,n,$
$x_{ik}\in\{0,1\},\ i=1,\cdots,n,\ k=1,\cdots,K,$
where $x_{ik}$ is represented in the same way as in (1). Note that the constraint $\sum\limits_{k}x_{ik}=1$, $x_{ik}\in\{0,1\},\ \forall k$, describes the semi-assignment constraint Integer-programming . Therefore, (2) is a semi-assignment optimization problem.
Here, $f(x)$ is a certain criterion to measure the qualification of graph clustering. Various choices of $f$ can be used, such as no34 ; GC-BM . In this paper, we consider the following form of objective function in GC-BM ,
$f(x)=\sum\limits_{i}\sum\limits_{j>i}\sum\limits_{k}x_{ik}x_{jk}d_{ij},$
where $d_{ij}$ represents the distance between the vertex $i$ and $j$. Below we mainly consider the following three choices of distances which are Burt’s distance Burt-distance ; jaccard-3 ; jaccard-4 , Jaccard distance jaccard-1 ; jaccard-3 ; jaccard-4 and Otsuka-Ochiai distance O-O-distance ; jaccard-3 ; jaccard-4 . As will be shown below, this objective function exhibits block properties.
To conclude this subsection, we would like to mention that there are also other criteria for $f$. For example, no34 focuses on maximizing the mean intra-cluster density while introducing a penalty function $P_{k}(M)$ to discourage clusters that are too large or too small. The penalty functions can be defined as $P_{k}(M)=\max\{0,\sum\limits_{i}x_{ik}\}$ or $P_{k}(M)=(\sum\limits_{i}x_{ik}-M)^{2}$, where $M$ is a parameter. Then, the objective function becomes:
$f(x)=-\sum\limits_{k=1}^{|\mathbb{C}|}\left[\sum\limits_{i,j}\left(\frac{w_{i,j}x_{ik}x_{jk}}{0.5\times n^{(k)}(n^{(k)}-1)}-\lambda P_{k}(M)\right)\right],$
where $w_{i,j}$ is the weight of the edge connecting vertex $i$ and vertex $j$, and $\lambda$ is a penalty coefficient. Unfortunately, this objective function does not exhibit block properties (Prop. 3) and is computationally challenging.
2.2 Different distance metrics
As mentioned above, graph clustering is defined as subsets of vertices that are considered similar to some extent. This similarity is manifested through the number of shared connections and is translated into a distance metric. Therefore, the distance measure we need is based on similarity rather than the shortest path distance. The following lists three types of distance that are often used as similarity measures in graph clustering problems.
Burt’s distance $D^{B}$
The Burt distance jaccard-3 ; jaccard-4 , borrowed from sociology Burt-distance , is defined as follows for the distance between the vertex $i$ and $j$:
$D^{B}_{ij}=\sqrt{\sum\limits_{s\neq i,j}(A_{is}-A_{js})^{2}},$
where $A=(A_{ij})\in\mathbb{R}^{n\times n}$ is the adjacency matrix. For an unweighted graph, $A_{ij}=1$ if $(i,j)\in E(G)$ and 00 otherwise. For a weighted graph, if $(i,j)\in E(G)$, $A_{ij}$ is the weight of the edge between $i$ and $j$; otherwise, $A_{ij}=0$. Then Burt’s distance matrix is defined as $D^{B}=(D^{B}_{ij})\in\mathbb{R}^{n\times n}$.
Jaccard distance $D^{J}$
The Jaccard distance jaccard-3 ; jaccard-4 , originating from botany jaccard-1 , is defined between the vertex $i$ and $j$ as follows:
$D^{J}_{ij}=\left\{\begin{array}[]{l}1-\frac{|a_{i}\cap a_{j}|}{|a_{i}\cup a_{j}|},\ i\neq j,\\ 0,\hskip 36.98866pti=j,\end{array}\right.$
where $a_{i}$ denotes the set of all vertices that share an edge with vertex $i$. The ratio $\frac{|a_{i}\cap a_{j}|}{|a_{i}\cup a_{j}|}$ represents the Jaccard similarity. The above applies to the case of an unweighted graph. For a weighted graph, we use the following expression:
$D^{J}_{ij}=\left\{\begin{array}[]{l}1-\frac{\sum\limits_{s=1}^{n}\min\{w_{is},w_{js}\}}{\sum\limits_{s=1}^{n}\max\{w_{is},w_{js}\}},\ i\neq j,\\ 0,\ i=j,\end{array}\right.$
where $w_{is}$ denotes the weight of the edge between the vertex $i$ and the vertex $s$. Note that $D^{J}_{ij}\in[0,1]$. Then Jaccard distance matrix is defined as $D^{J}=(D^{J}_{ij})\in\mathbb{R}^{n\times n}$.
Otsuka-Ochiai distance $D^{O}$
The Otsuka-Ochiai distance jaccard-3 ; jaccard-4 originates from zoology O-O-distance . In the case of unweighted gaphs, Otsuka-Ochiai distance between the vertex $i$ and $j$ is defined as follows:
$D^{O}_{ij}=1-\frac{|a_{i}\cap a_{j}|}{\sqrt{|a_{i}|\times|a_{j}|}}\in[0,1],$
where $a_{i}$ is the same as that in the Jaccard distance. In the weighted case, we use the following expression:
$D^{O}_{ij}=1-\frac{\sum\limits_{s=1}^{n}\min\{w_{is},w_{js}\}}{\sqrt{\sum\limits_{s=1}^{n}w_{is}\times\sum\limits_{s=1}^{n}w_{js}}}\in[0,1],$
where $w_{is}$ denotes the weight of the edge between the vertex $i$ and $s$. Then Otsuka-Ochiai distance matrix is defined as $D^{O}=(D^{O}_{ij})\in\mathbb{R}^{n\times n}$.
The three $n\times n$ distance matrices (denoted as $D^{B}$, $D^{J}$ and $D^{O}$ respectively) satisfy Prop. 1.
Proposition 1
For the matrix $D$ defined by $D^{B}$, $D^{J}$ or $D^{O}$ as above, it holds that $D_{ij}\geqslant 0$, $D_{ii}=0$, $D_{ij}=D_{ji}$, $i=1,\cdots,n$, $j=1,\cdots,n$. In other words, $D$ is nonnegative, symmetric and has zero diagonal elements.
[…]
Based on Sect. 2.1 and Sect. 2.2, the graph clustering model we consider in this paper takes the following form
$\min\limits_{x}$ $\sum\limits_{i}\sum\limits_{j>i}\sum\limits_{k}x_{ik}x_{jk}d_{ij}$ (3)
$s.t.$ $\sum\limits_{k}x_{ik}=1,\ i=1,\cdots,n,$
$x_{ik}\in\{0,1\},\ i=1,\cdots,n,\ k=1,\cdots,K.$
Denote ${\bf x}=({\bf x}^{\top}_{1},{\bf x}^{\top}_{2},\cdots,{\bf x}^{\top}_{n})^{\top}\in\mathbb{R}^{nK}$, ${\bf x}^{\top}_{i}=(x_{i1},x_{i2},\cdots,x_{iK})\in\mathbb{R}^{K}$. Then the sum of distances generated by the vertex $i$ and $j$ can be expressed as the following equation
$\sum\limits_{k}x_{ik}x_{jk}d_{ij}=\left[\begin{array}[]{c c c c}x_{i1}\ x_{i2}\ \cdots\ x_{iK}\end{array}\right]\left[\begin{array}[]{c c c c}d_{ij}&0&\cdots&0\\ 0&d_{ij}&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&d_{ij}\end{array}\right]\left[\begin{array}[]{c}x_{j1}\\ x_{j2}\\ \vdots\\ x_{jK}\end{array}\right]={\bf x}_{i}^{\top}(d_{ij}I_{k}){\bf x}_{j},$ (4)
where $I_{k}\in\mathbb{R}^{K\times K}$ is an identity matrix. Let $\overline{\Delta}_{ij}=d_{ij}I_{k}\in\mathbb{R}^{K\times K}$ and $\Delta=D\otimes I_{k}\in\mathbb{R}^{nK\times nK}$. Then the objective function of (3) can be written as
$\sum\limits_{i}\sum\limits_{j>i}\sum\limits_{k}x_{ik}x_{jk}d_{ij}=\frac{1}{2}\left[\begin{array}[]{c c c c}{\bf x}_{1}^{\top}\ {\bf x}_{2}^{\top}\ \cdots\ {\bf x}_{n}^{\top}\end{array}\right]\left(D\otimes I_{k}\right)\left[\begin{array}[]{c}{\bf x}_{1}\\ {\bf x}_{2}\\ \vdots\\ {\bf x}_{n}\end{array}\right]=\frac{1}{2}{\bf x}^{\top}\Delta{\bf x}.$ (5)
Hence, (3) can be written in the following equivalent form
$\min\limits_{{\bf x}}$ $\hskip 5.69054ptf({\bf x}):=\frac{1}{2}{\bf x}^{\top}\Delta{\bf x}$ (6)
${\rm s.t.}$ $\hskip 5.69054pt{\bf x}_{i}^{\top}{\bf 1}_{k}=1,\ i=1,\cdots,n,$
$\hskip 5.69054pt\textbf{x}\in\{0,1\}^{nK},$
where ${\bf 1}_{k}=[1,1,\cdots,1]^{\top}\in\mathbb{R}^{K}$. (6) can also be written in the following equivalent form:
$\min\limits_{{\bf x}\in\mathbb{R}^{nK}}$ $\hskip 5.69054ptf({\bf x})$ (7)
${\rm s.t.}$ $\hskip 5.69054pt{\bf x}_{i}^{\top}{\bf 1}_{k}=1,\ i=1,\cdots,n,$
$\hskip 5.69054pt{\bf x}\geqslant 0,$
$\hskip 5.69054pt\|{\bf x}\|_{0}\leqslant n.$
In (7), we remove the last constraint ${\bf x}\in\{0,1\}^{nK}$ of (6), and add two constraints, $\|{\bf x}\|_{0}\leqslant n$ and ${\bf x}\geqslant 0$. Due to the constraint ${\bf x}_{i}^{\top}{\bf 1}_{k}=1$, $i=1,\cdots,n$, we know that $\|{\bf x}\|_{0}\geqslant n$. Additionally, because of the last constraint $\|{\bf x}\|_{0}\leqslant n$, we have $\|{\bf x}\|_{0}=n$.
The following result is obvious due to the definition of $\Delta$.
Proposition 2
Recall $\Delta=D\otimes I_{k}$. It holds that $\Delta_{ii}=0$, $\Delta_{ij}=\Delta_{ji}$, $i=1,2,\cdots,nK$, $j=1,2,\cdots,nK$. In other words, $\Delta$ is nonnegative, symmetric and has zero diagonal elements.
To present the property of $f(x)$, let an index $i_{0}\in\{1,2,\cdots,n\}$ and the set $I^{-i_{0}}=\{1,2,\cdots,i_{0}-1,i_{0}+1,\cdots,n\}$. We can rewrite $f({\bf x})$ in (6) as:
$f({\bf x})$ $=\frac{1}{2}\sum\limits_{i=1}^{n}\sum\limits_{j\neq i}{\bf x}_{i}^{\top}\overline{\Delta}_{ij}{\bf x}_{j}$ (8)
$=\frac{1}{2}\left(2{\bf x}_{i_{0}}^{\top}\sum\limits_{j\neq i_{0}}\overline{\Delta}_{i_{0}j}{\bf x}_{j}+\sum\limits_{i\in I^{-i_{0}}}\sum\limits_{j\neq i,j\in I^{-i_{0}}}{\bf x}_{i}^{\top}\overline{\Delta}_{ij}{\bf x}_{j}\right)$
$={\bf x}_{i_{0}}^{\top}\sum\limits_{j\neq i_{0}}\overline{\Delta}_{i_{0}j}{\bf x}_{j}+\frac{1}{2}\sum\limits_{i\in I^{-i_{0}}}\sum\limits_{j\neq i,j\in I^{-i_{0}}}{\bf x}_{i}^{\top}\overline{\Delta}_{ij}{\bf x}_{j}$
$:=f^{(i_{0})}({\bf x})+f^{(-i_{0})}({\bf x}).$
We have the following result.
Proposition 3
$f^{(i)}({\bf x})={\bf x}_{i}^{\top}\nabla_{{\bf x}_{i}}f({\bf x})$, $i\in\{1,2,\cdots,n\}$.
Proof. According to (4) and (5), we know that
$f({\bf x})=\frac{1}{2}{\bf x}^{\top}\Delta{\bf x}=\sum\limits_{i}\sum\limits_{j>i}{\bf x}_{i}^{\top}\overline{\Delta}_{ij}{\bf x}_{j}.$
Then $\nabla_{{\bf x}_{i}}f({\bf x})=\sum\limits_{j\neq i}^{n}\overline{\Delta}_{ij}{\bf x}_{j}$. Due to (8), we have
$f^{(i_{0})}({\bf x})={\bf x}_{i_{0}}^{\top}\sum\limits_{j\neq i_{0}}\overline{\Delta}_{i_{0}j}{\bf x}_{j}={\bf x}_{i_{0}}^{\top}\nabla_{{\bf x}_{i_{0}}}f({\bf x}).$
Then for every index $i\in\{1,2,\cdots,n\}$, $f^{(i)}({\bf x})={\bf x}_{i}^{\top}\nabla_{{\bf x}_{i}}f({\bf x}).$Box
Remark 1
The above proposition demonstrates that for each block ${\bf x}_{i}$, $f({\bf x})$ is basically a linear function with respect to ${\bf x}_{i}$. This is a key property which will be further explored in the subsequent of the paper.
[…]
In this section, we relax the sparse constraint problem (7) to a continuous problem and propose a fast quadratic penalty algorithm to solve the relaxation problem.
3.1 Relaxation problem
Both (6) and (7) are essentially a discrete optimization problem, which is in general NP hard and therefore is extremely difficult to solve. A popular way to deal with (6) or (7) is to relax the discrete constraint and consider solving the relaxed continuous problem. By removing the last constraint in (7), we obtain the following relaxation problem
$\min\limits_{{\bf x}\in\mathbb{R}^{nK}}$ $\hskip 5.69054ptf({\bf x})$ (9)
${\rm s.t.}$ $\hskip 5.69054pt{\bf x}_{i}^{\top}{\bf 1}_{k}=1,\ i=1,\cdots,n,$
$\hskip 5.69054pt{\bf x}\geqslant 0.$
(9) is a continuous problem with simplex constraints. Due to Prop. 3, the following result holds by Thm. 1 in no37 , which address the relation between (7) and the relaxation problem (9).
Theorem 3.1
There exists a global minimum ${\bf x}^{*}$ of the problem (9) such that $\|{\bf x}^{*}\|_{0}=n$, and the global minimizer ${\bf x}^{*}\in\mathbb{R}^{nK}$ is also a global minimizer of the problem (7).
Based on Thm. 3.1, starting from any global minimizer of problem (9), we can eventually find a point ${\bf x}^{*}$ that is a global minimizer of both (9) and (7). Therefore, we can find a global minimizer of (7) using the approach described in Alg. 1 below.
Algorithm 1 The procedure of finding a global minimizer of (7)
1:  Input: a global minimizer of (9): ${\bf y}^{0}=(({\bf y}^{0}_{1})^{\top},({\bf y}^{0}_{2})^{\top},\cdots,({\bf y}^{0}_{n})^{\top})^{\top}\in\mathbb{R}^{nK}$. Let ${\bf x}=(({\bf x}_{1})^{\top},({\bf x}_{2})^{\top},\cdots,({\bf x}_{n})^{\top})^{\top}={\bf 0}\in\mathbb{R}^{nK}$;
2:  for $i=1:n$ do
3:     For $i$-th block ${\bf y}^{0}_{i}$ of $y$, we find an index $p^{i}$ which $({\bf y}^{0}_{i})_{p^{i}}\geqslant({\bf y}^{0}_{i})_{q}$, $q=1,2,\cdots,K$.
4:     Let $({\bf x}_{i})_{p^{i}}=1$.
5:  end for
6:  Output: ${\bf x}=(({\bf x}_{1})^{\top},({\bf x}_{2})^{\top},\cdots,({\bf x}_{n})^{\top})^{\top}\in\mathbb{R}^{nK}$, which is a global minimizer of (7).
3.2 Quadratic penalty method
To solve the continuous relaxation problem (9), various optimization methods can be used. It has been verified in no36 ; no37 that the quadratic penalty method is highly efficient in solving such kind of problem. Therefore, in this paper, we continue to apply the quadratic penalty method to solve (9). The idea of this method is as follows. Due to the result in Thm. 3.1, there is no need to solve (9) to get an accurate global minimizer. All we need is to identify the support set of the global minimizer of (9) so that we can use Alg. 1 to obtain a global minimizer of (7). Therefore, we can penalize the equality constraint and in each iteration, we solve the quadratic penalty problem
$\min\limits_{0\leqslant{\bf x}\leqslant M}g^{l}({\bf x}):=f({\bf x})+\frac{\theta_{l}}{2}\sum\limits_{i=1}^{n}({\bf x}_{i}^{\top}{\bf 1}_{k}-1)^{2},$ (10)
where $M\geqslant 1$ is a given value and the upper bound ${\bf x}\leqslant M$ is added to make sure that the penalty problem is well-defined. Details of the quadratic penalty method are given in Alg. 2.
Algorithm 2 Quadratic penalty method
1:  Input: an initial point ${\bf x}^{0}$, a parameter $\theta_{0}>0$, $l=0$;
2:  for $l$ do
3:     Solve (10) to get a ${\bf x}^{l}$.
4:     if the termination rule is not satisfied then
5:        Choose parameter $\theta_{l+1}\geqslant\theta_{l}$; $l=l+1$;
6:     else
7:        break;
8:     end if
9:  end for
10:  Apply Alg. 1 to ${\bf x}^{l}$ to get a global minimizer of (7).
Theorem 3.2 (nno5 ; nno6 )
Let $\{{\bf x}^{l}\}$ be the sequence generated by Alg. 2, and assume that ${\bf x}^{l}$ is a global minimizer of (10). Let $\lim\limits_{l\rightarrow\infty}\theta_{l}=+\infty$. Then any accumulation point of this generated sequence is a global minimum of (9).
Thm. 3.2 addresses the convergence of the quadratic penalty method. A detailed proof can be found in nno5 ; nno6 .
Assumption 1
Let $\{{\bf x}^{l}\}$ be the sequence generated by Alg. 2, with $\lim\limits_{l\rightarrow\infty}\theta_{l}=+\infty$. Here, $l$ is a positive integer. Suppose $\lim\limits_{l\rightarrow\infty}{\bf x}^{l}={\bf z}$, and ${\bf z}$ is a global minimizer of (9).
Define the support set at ${\bf x}$ by $\Gamma({\bf x}):=\{j:\ {\bf x}_{j}>0\}$. Based on Thm. 3 and Thm. 4 in no37 , the following result holds.
Theorem 3.3
Under Assump. 1, there exists a global minimizer ${\bf z}^{*}$ of (7) such that for $l$ sufficiently large, it holds that $\Gamma({\bf x}^{l})=\Gamma({\bf z}^{*})$.
Thm. 3.3 shows that under Assump. 1, the support set of the global minimizer for the original problem (7) can be precisely recovered when the number of iterations is sufficiently large.
Inspired by no37 , we employ a projected gradient method based on the active set to solve the subproblem (10), as detailed in no37 Alg. 3. Our primary goal is to identify the support set of the global minimizer of (7), rather than its magnitude. Therefore, the projected gradient method is applied to the nonlinear problem (10) with simple box constraints.
4 Numerical Results
In this section, we evaluate the performance of our algorithm on different data sets. We compare the solution quality of our algorithm (QP-GC) with the leading commercial solver Gurobi Gurobi for (3) and the Boltzmann machine (BM) from the latest research GC-BM . Additionally, we conduct case studies on two real-world graphs. We implement the Boltzmann machine by ourselves, since the original code is not available. Our algorithm, the Boltzmann machine and Gurobi run on a 4-core/8-thread machine. All experiments are written in MATLAB R2022a [^1] https://ww2.mathworks.cn running in Windows 10 on an Intel(R) Core(TM) i7-1065G7 CPU at 1.30 GHZ with 16.0 GB of RAM, and all graphs are generated using Python 3.8222https://www.python.org/downloads/.
4.1 Evaluation of clustering quality
To start with, we introduce some measures to evaluate the clustering quality. We use the comparison of intra-cluster density, inter-cluster density, and overall density as metrics to evaluate clustering quality. It is demonstrated in cq-1 and cq-2 that the evaluation of the quality of the cluster using these metrics is far superior to the most popular graph cluster quality function, i.e., modularity cq-3 .
The number of edges connecting vertices within cluster $i$ is denoted as $|E_{ii}|$, and the number of edges connecting a vertex in cluster $i$ to a vertex in cluster $j$ is denoted as $|E_{ij}|$. The overall density is defined by
$\kappa=\frac{|E(G)|}{0.5\times n(n-1)}.$
For a cluster $i$, the intra-cluster density is given by
$\kappa^{(i)}_{intra}=\frac{|E_{ii}|}{0.5\times n^{(i)}(n^{(i)}-1)}.$
For a cluster $i$ and a cluster $j$, $i\neq j$, the inter-cluster density is given by
$\kappa^{(ij)}_{inter}=\frac{|E_{ij}|}{n^{(i)}\times n^{(j)}}.$
For a graph clustered into $K$ clusters, the mean intra-cluster density and the mean inter-cluster density are given respectively by
$\bar{\kappa}_{intra}=\frac{1}{K}\sum\limits^{K}_{i=1}\kappa^{(i)}_{intra},\ \bar{\kappa}_{inter}=\frac{1}{0.5\times K(K-1)}\sum\limits^{K}_{i=1}\sum\limits^{K}_{j=i+1}\kappa^{(ij)}_{inter}.$
Using the quantities defined above, we can measure the quality of clustering. A high quality clustering groups vertices into clusters such that, on average, the links between vertices within these clusters are denser than the links between vertices in different clusters cq-1 ; cq-2 . Therefore, a high quality clustering should satisfy the following inequality
$\bar{\kappa}_{inter}<\kappa<\bar{\kappa}_{intra}.$ (11)
For the choice of distance, considering factors such as wide applicability, computational complexity, and interpretability, we take the Jaccard distance as the measurement to report our result. Furthermore, it is both demonstrate in jaccard-3 ; jaccard-4 that the Jaccard distance is the most suitable distance metric for graph clustering.
4.2 Synthetic graphs
In this part, we perform experiments on nine different synthetic graphs, generated using two different models: the Planted Partition Model (PPM) PPM and the Stochastic Block Model (SBM) SBM . These graphs have known classification properties, as detailed in Tab. 1. The graphs are generated using the NetworkX Python library no26 . The clustering process becomes more complicated as intra-cluster edge probability decreases, inter-cluster edge probability increases, and cluster sizes vary. An increase in the number of graph vertices also adds to the computational difficulty.
It is worth noting that the mean intra-cluster density $\bar{\kappa}_{intra}$ (or the mean inter-cluster density $\bar{\kappa}_{inter}$) is an empirical estimate of the intra-cluster edge probability ${\rm P_{intra}}$ (or the inter-cluster edge probability ${\rm P_{inter}}$) in the generated models, as has been demonstrated by GC-BM . Therefore, the closer $\bar{\kappa}_{intra}$ (or $\bar{\kappa}_{inter}$) calculated from the clustering obtained by a given method is to ${\rm P_{intra}}$ (or ${\rm P_{inter}}$) of the generation model, the higher the quality of the solution generated by that method. This also indicates that the clustering performance of the method is better.
Table 1: Details of nine synthetic graphs.
Model type Graph name ${\rm P_{intra}}$ ${\rm P_{inter}}$ $K$ $n$ Cluster sizes
PPM G1_PPM 0.9 0.1 5 250 50
PPM G2_PPM 0.85 0.15 5 250 50
PPM G3_PPM 0.8 0.2 5 250 50
SBM G1_SBM 0.9 0.05 30 3000 [25,200]
SBM G2_SBM 0.9 0.1 30 3000 [25,200]
SBM G3_SBM 0.85 0.05 30 3000 [25,200]
SBM G4_SBM 0.85 0.1 30 3000 [25,200]
SBM G5_SBM 0.8 0.05 30 3000 [25,200]
SBM G6_SBM 0.8 0.1 30 3000 [25,200]
4.2.1 Performance for QP-GC
We first test on three PPM graph models with $250$ vertices, each containing five clusters with $50$ vertices per cluster. Secondly, we test on six SBM graph models with $3000$ vertices. Each graph contains $30$ clusters, with cluster sizes ranging from $25$ to $200$ vertices. For each intra-cluster edge probability, two different inter-cluster edge probabilities are used to create the graphs. These graphs are generated using the following intra-cluster/inter-cluster edge probabilities (${\rm P_{intra}}$/${\rm P_{inter}}$) as shown in Tab. 1. The variability in cluster sizes, along with the large number of vertices, adds the complexity to the clustering problem.
The proposed QP-GC method for solving graph clustering problems is tailored to the problem (3). This is due to the objective function and its gradient given by $\nabla_{{\bf x}_{i}}f({\bf x})=\sum_{j\neq i}\overline{\Delta}_{ij}{\bf x}_{j}$. Additionally, (3) involves only a single vector variable ${\bf x}>0$ and $n$ linear equality constraints, all of which are linear. These characteristics ensure that the computational process runs efficiently.
The parameters for our method are set as follows. In Alg. 2, we set the initial point ${\bf x}^{0}$ as a random matrix. Update $\theta_{l}$ as
$\theta_{l+1}=\left\{\begin{array}[]{ll}3\theta_{l},&{\rm if}\ \|h^{l}\|>0.01\ {\rm and}\ \theta_{l}<\bar{\theta},\\ \theta_{l},&{\rm otherwise},\end{array}\right.$
where $h^{l}=\left(({\bf x}^{l}_{1})^{\top}{\bf 1}-1,\cdots,({\bf x}^{l}_{n})^{\top}{\bf 1}-1\right)$. For the experiments on the PPM model, we set $\bar{\theta}=10^{5}$ and $\theta_{0}=10$, while for the experiments on the SBM model, we set $\bar{\theta}=10^{10}$ and $\theta_{0}=2500$. Each ${\bf x}^{l}$ returned is projected onto a binary assignment matrix using Alg. 1. For the experiments on the small-scale PPM model, we set $\epsilon=0.0009$, while for the large-scale SBM model experiments, we set $\epsilon=2\times 10^{-8}$.
Results on PPM graph models
We test on three PPM models. In Tab. 2, $\epsilon_{intra}$ (or $\epsilon_{inter}$) represents the absolute value of the difference between $\bar{\kappa}_{intra}$ (or $\bar{\kappa}_{inter}$) computed by QP-GC and ${\rm P_{intra}}$ (or ${\rm P_{inter}}$) in the graph. We say that the smaller its value, the more accurate the clustering result, especially with regard to $\epsilon_{intra}$. We also report the CPU time and the number of iterations $C_{iter}$ for QP-GC.
Table 2: Numerical results of QP-GC for three PPM models with $n=250$ and $K=5$.
Graph name Graph characteristics QP-GC
${\rm P_{intra}}$ ${\rm P_{inter}}$ $\kappa$ $\bar{\kappa}_{intra}$ $\bar{\kappa}_{inter}$ $\epsilon_{intra}$ $\epsilon_{inter}$ times(s) $C_{iter}$
G1_PPM 0.9 0.1 0.26 0.90 0.10 0 0 0.6 230
G2_PPM 0.85 0.15 0.29 0.85 0.15 0 0 0.8 320
G3_PPM 0.8 0.2 0.32 0.80 0.20 0 0 0.9 430
To provide a more intuitive demonstration of QP-GC, we visualized the above three graphs before and after clustering, as shown in Fig. 2. The visualizations of the three PPM graphs are shown in the left column of Fig. 2. The right column of Fig. 2 show the graph after clustering with QP-GC, where the graph is divided into five clusters, with each color representing a different cluster. Fig. 2 illustrates that our method accurately partitions three graphs into five clusters, demonstrating high-quality clustering. This visual representation is sufficient to show the effectiveness of QP-GC.
(a) The original graph G1_PPM
(b) Graph G1_PPM after clustering
(c) The original graph G2_PPM
(d) Graph G2_PPM after clustering
(e) The original graph G3_PPM
(f) Graph G3_PPM after clustering
Figure 2: A visual representation of three PPM graph models after clustering with our method
Results on SBM graph models
We test on six SBM models. In Tab. 3, $\epsilon_{intra}$ (or $\epsilon_{inter}$) is defined in the same way as in Tab. 2. Additionally, we also list the CPU time and the number of iterations for our method. Note that the experiments are conducted on SBM models with $3000$ nodes, $30$ clusters, and the number of nodes in each cluster randomly chosen from the range $[25,200]$. Due to the complexity of this scenario, visualizations, as in Fig. 2, are not intuitive. However, $\epsilon_{intra}$ (or $\epsilon_{inter}$) is sufficient to demonstrate the clustering performance. As shown in Tab. 3, the more complex the graph, the greater the number of iterations required and the longer the CPU time.
Table 3: Numerical results of QP-GC for six SBM models with $n=3000$ and $K=30$.
Graph name Graph characteristics QP-GC
${\rm P_{intra}}$ ${\rm P_{inter}}$ $\kappa$ $\bar{\kappa}_{intra}$ $\bar{\kappa}_{inter}$ $\epsilon_{intra}$ $\epsilon_{inter}$ times(mm:ss) $C_{iter}$
G1_SBM 0.90 0.05 0.09 0.78 0.06 0.12 0.01 17:07 2930
G2_SBM 0.90 0.10 0.14 0.73 0.12 0.17 0.02 43:17 6000
G3_SBM 0.85 0.05 0.09 0.73 0.06 0.12 0.01 27:26 4560
G4_SBM 0.85 0.10 0.13 0.72 0.11 0.13 0.01 43:39 6630
G5_SBM 0.80 0.05 0.08 0.65 0.06 0.15 0.01 30:16 5000
G6_SBM 0.80 0.10 0.13 0.66 0.11 0.14 0.01 44:33 6750
4.2.2 Comparison with other algorithms
Results on PPM graph models
The numerical results are presented in Tab. 4. We compare QP-GC with Gurobi and BM in terms of $\epsilon_{intra}$, $\epsilon_{inter}$, and CPU time. From Tab. 5, we can observe that all three solutions produce the same results, and all satisfy (11). In all three cases (G1_PPM, G2_PPM and G3_PPM), they perfectly recover the generation model. We would like to emphasize that in this example, it can be noticed that under the same conditions, our method achieves accurate clustering results in a shorter CPU time.
Table 4: Numerical results of three methods for three PPM models with $n=250$ and $K=5$.
Graph name Graph characteristics QP-GC Gurobi BM
${\rm P_{intra}}$ ${\rm P_{inter}}$ $\kappa$ $\bar{\kappa}_{intra}$ $\bar{\kappa}_{inter}$ $\bar{\kappa}_{intra}$ $\bar{\kappa}_{inter}$ $\bar{\kappa}_{intra}$ $\bar{\kappa}_{inter}$
G1_PPM 0.90 0.10 0.26 0.90 0.10 0.90 0.10 0.90 0.10
G2_PPM 0.85 0.15 0.29 0.85 0.15 0.85 0.15 0.85 0.15
G3_PPM 0.80 0.20 0.32 0.80 0.20 0.80 0.20 0.80 0.20
Table 5: Comparison of some clustering metrics on PPM models.
Methods G1_PPM G2_PPM G3_PPM
$\epsilon_{intra}$ $\epsilon_{inter}$ times(s) $\epsilon_{intra}$ $\epsilon_{inter}$ times(s) $\epsilon_{intra}$ $\epsilon_{inter}$ times(s)
QP-GC 0 0 0.6 0 0 0.8 0 0 0.9
Gurobi 0 0 3.2 0 0 3.1 0 0 3.1
BM 0 0 3.5 0 0 3.6 0 0 3.7
Results on SBM graph models
The characteristics of the graphs and the numerical results are reported in Tab. 6. Similarly, we compare QP-GC with Gurobi and BM in terms of $\epsilon_{intra}$, $\epsilon_{inter}$, and CPU time. Using the CPU time required by QP-GC to obtain the results as a benchmark, we compare the clustering quality returned by the three methods within that time period. Based on the results in Tab. 6 and Tab. 7, we summarize the following two observations:
  • (a)
    It is observed that the results obtained by QP-GC and BM both satisfy the inequality (11). However, Gurobi fails to return meaningful results under the same conditions, and the inequality does not hold (indicated by underline in Tab. 6).
  • (b)
    We notice significant differences in the results among the solvers. As shown in Tab. 7, within the same CPU time, our method achieves better clustering quality for this type of model compared to the solution returned by BM. In the case of larger and more complex SBM graphs, our method provides both faster and more reliable results for these models.
Table 6: Numerical results of three methods for six SBM models with $n=3000$ and $K=30$.
Graph name Graph characteristics QP-GC Gurobi BM
${\rm P_{intra}}$ ${\rm P_{inter}}$ $\kappa$ $\bar{\kappa}_{intra}$ $\bar{\kappa}_{inter}$ $\bar{\kappa}_{intra}$ $\bar{\kappa}_{inter}$ $\bar{\kappa}_{intra}$ $\bar{\kappa}_{inter}$
G1_SBM 0.90 0.05 0.09 0.78 0.06 0.33 0.32 0.10 0.09
G2_SBM 0.90 0.10 0.14 0.72 0.12 0.32 0.32 0.15 0.14
G3_SBM 0.85 0.05 0.09 0.73 0.06 0.30 0.29 0.10 0.08
G4_SBM 0.85 0.10 0.13 0.72 0.11 0.34 0.34 0.14 0.13
G5_SBM 0.80 0.05 0.08 0.65 0.06 0.26 0.26 0.10 0.08
G6_SBM 0.80 0.10 0.13 0.66 0.11 0.23 0.24 0.14 0.13
Table 7: Comparison of some clustering metrics on SBM models.
Methods G1_SBM  (times  17:07) G2_SBM  (times  43:17) G3_SBM  (times  27:26)
$\epsilon_{intra}$ $\epsilon_{inter}$ $\epsilon_{intra}$ $\epsilon_{inter}$ $\epsilon_{intra}$ $\epsilon_{inter}$
QP-GC 0.12 0.01 0.17 0.02 0.12 0.01
Gurobi 0.57 0.27 0.58 0.22 0.55 0.24
BM 0.80 0.04 0.75 0.04 0.75 0.03
Methods G4_SBM  (times  43:39) G5_SBM  (times  30:16) G6_SBM  (times  44:33)
$\epsilon_{intra}$ $\epsilon_{inter}$ $\epsilon_{intra}$ $\epsilon_{inter}$ $\epsilon_{intra}$ $\epsilon_{inter}$
QP-GC 0.13 0.01 0.15 0.01 0.14 0.01
Gurobi 0.51 0.24 0.54 0.21 0.57 0.14
BM 0.71 0.03 0.70 0.03 0.66 0.03
To clearly and intuitively demonstrate the superiority of our method, we represent ${\rm P_{intra}}$ and $\bar{\kappa}_{intra}$ from Tab. 6 using radar plots, as shown in Fig. 3(a), and similarly represent ${\rm P_{inter}}$ and $\bar{\kappa}_{inter}$ using radar plots, as shown in Fig. 3(b). From Fig. 3, we observe that, regardless of the SBM graph model, $\bar{\kappa}_{intra}$ and $\bar{\kappa}_{inter}$ produced by our method (represented by the blue line) are closer to the ${\rm P_{intra}}$ and ${\rm P_{inter}}$ of the graph model (represented by the red line).
(a) Comparison of ${\rm P_{intra}}$ and $\bar{\kappa}_{intra}$
(b) Comparison of ${\rm P_{inter}}$ and $\bar{\kappa}_{inter}$
Figure 3: The radar plot representation of ${\rm P_{intra}}$ (or ${\rm P_{inter}}$) and $\bar{\kappa}_{intra}$ (or $\bar{\kappa}_{inter}$) in Tab. 6
4.3 Real-world graphs
Although synthetic networks provide a repeatable and controlled testing platform for our experiments, testing algorithms on real-world network data is also necessary, even if real-world graphs are often not the best benchmarks for assessing clustering quality. In this section, we select two representative real-world network datasets: Zachary’s Karate Club (Karate_club) no3 and the United States College Football Division IA 2000 season graph (US_football_2000) no3 . Both contain a known clustering structure. The graph data used in our experiments are available at https://github.com/qinyuenlp/CommunityDetection/tree/master/data. The characteristics of these graphs are shown in Tab. 8.
Table 8: Graph characteristics of two real-world graphs.
Graph name Graph characteristics
$|V|$ $|E|$ $\kappa$ Clusters
Karate_club 34 78 0.14 2
US_football_2000 115 613 0.09 12
For experiments on real-world graphs, the following two points must be noted: (1) Real-world graphs are often instances of unknown underlying generative models with random noise. In many cases, modifying the clustering assignment of real-world graph vertices may result in higher intra-cluster density. This means that the true clustering results of the real-world graph vertices may not correspond to the clusters with the highest density. (2) When comparing the clustering results with the true clusters in real-world graphs, it is not sufficient to simply compare the cluster labels. Instead, the comparison should focus on the composition of the clusters. For instance, the cluster $C_{1}$ in the true clustering might not correspond to $C_{1}$ from the QP-GC clustering result, but it might fully correspond to $C_{2}$ from the QP-GC result. These two factors determine that we need to set evaluation criteria different from those used in the synthetic graph experiments.
For both of the following experiments, we use the same parameter settings. $\theta_{0}=20$ and the initial point $S^{0}$ is set as a random matrix. We set $\epsilon=10^{-5}$ in no37 Alg. 3. For the termination condition, we set $\tau=0.01$. Other parameters are the same as the PPM experiments.
4.3.1 Karate club study of Zachary
The real-world graph network data in this subsection comes from Zachary’s famous study of the karate club no3 . The vertices of the graph represent the $34$ members of the karate club, and the edges represent the social connections between the members. Here, we use the simplified unweighted version of the network. The club split into two factions due to internal conflict, and this division is reflected in the graph structure, as shown in Fig. 4(a). Note that in each subfigure of Fig. 4, nodes with the same color indicate that they belong to the same cluster. However, the same color across different subfigures does not convey any particular meaning.
(a) The true clustering
(b) The clustering result obtained by QP-GC
(c) The clustering result from Gurobi
(d) The clustering result from BM
Figure 4: The friendship network derived from Zachary’s karate club, as described in this paper. Nodes of different colors represent membership in different factions. The red squares in the figure indicate the nodes of clustering error
We apply our method to this graph to identify the factions involved in the club’s split. In addition, under the same CPU time ($1$ second), we compare the clustering results obtained using Gurobi and BM with those produced by our method. Both Gurobi and BM have been shown to be effective for clustering in GC-BM . Since the graph has two cluster structures, we visualize the clustering, making the results visually clear, as shown in Fig. 4. We find that our QP-GC method misclassifies only one node, as shown in Fig. 4(b), while Gurobi misclassifies $11$ nodes, as shown in Fig. 4(c), and BM misclassifies two nodes, as shown in Fig. 4(d). Thus, our method more accurately restores the clustering. Specifically, we observe six nodes circled by the red dashed ellipse in Fig. 4. These six nodes are shown as belonging to the same cluster in Fig. 4(a), and in Fig. 4(b), they are classified into the same cluster by QP-GC. However, in Fig. 4(c) and Fig. 4(d), it is shown that these nodes are not grouped together.
4.3.2 US College Football Division IA 2000 season graph
This subsection uses the well-known United States College Football Division IA $2000$ season graph no3 , which represents the matchups between $115$ college football teams during the regular season. The vertices in the graph represent the teams, labeled by their university names, and the edges indicate that the two connected teams have faced each other at least once during the regular season. These teams are divided into $12$ conferences, with each conference containing approximately $8$-$12$ teams. Teams within the same conference compete more frequently against each other than against teams from different conferences, resulting in more shared connections among teams in the same conference compared to those in different ones.
We also compared the clustering results of QP-GC with Gurobi and BM when running for the same CPU time ($1$ second). The graph consists of $12$ clusters, as shown in Fig. 5, where each circle represents a cluster. We assign $12$ different colors to these circles. QP-GC misclassifies $12$ nodes, as shown in Fig. 5(b), while Gurobi misclassifies $60$ nodes, as shown in Fig. 5(c) and BM misclassifies $17$ nodes, as shown in Fig. 5(d). Therefore, our method more accurately recovers the clustering structure. Specifically, we observe the $11$ nodes circled by the red dashed ellipse in Fig. 5. These $11$ nodes originally belong to the same cluster (as shown in Fig. 5(a)). In Fig. 5(b), QP-GC classifies these $11$ nodes into the same cluster, while in Fig. 5(c) and Fig. 5(d), it is clear that they are not grouped together.
(a) The true clustering
(b) The clustering result obtained by QP-GC
(c) The clustering result from Gurobi
(d) The clustering result from BM
Figure 5: A visualization of the US College Football Division IA 2000 season graph. Each circle represents nodes (teams) belonging to the same cluster, while red squares indicate misclassified nodes
The large number of clusters ($12$ cluster structures) makes Fig. 5 less intuitive and clear. To better compare the clustering results, and considering the two key factors mentioned earlier for experiments on real-world graphs, we also use the Jaccard similarity function $J_{C^{\psi}_{j}}$ to compare the compositions of the true clusters and the clusters identified by various methods.
This is expressed as
$J_{C^{\psi}_{j}}=\max\limits_{C^{true}_{i}}\left\{\frac{|C^{true}_{i}\cap C^{\psi}_{j}|}{|C^{true}_{i}\cup C^{\psi}_{j}|}\right\},$
where $C^{true}_{i}$ represents the set of nodes in the $i$-th true cluster, and $C^{\psi}_{j}$ denotes the set of nodes in the $j$-th cluster obtained by method $\psi$. We know $0\leqslant J_{C^{\psi}_{j}}\leqslant 1$. The value produced by a perfect match between the $j$-th cluster identified by method $\psi$ and its corresponding true cluster is $J_{C^{\psi}_{j}}=1$, while a complete mismatch results in $J_{C^{\psi}_{j}}=0$. We then have a set of $J_{C^{\psi}_{j}}$ defined as $\mathbb{J}^{\psi}=\{J_{C^{\psi}_{j}},\ j=1,\cdots,k\}$. Let $\mathbb{J}^{\psi}_{p}=\{J_{C^{\psi}_{j}}:\ J_{C^{\psi}_{j}}=1,\ j=1,\cdots,k\}$, we calculate $|\mathbb{J}^{\psi}_{p}|$. We also calculate the mean value $\bar{\mathbb{J}}^{\psi}$ of all elements in $\mathbb{J}^{\psi}$ to provide an intuitive and effective comparison of the results, where a larger mean indicates better clustering performance. The comparison is listed in Tab. 9. These data show that our method achieves a better match to the true clusters for this class of optimization models.
Table 9: Some reference values related to $J_{C^{\psi}_{j}}$ for comparison with true clusters.
Reference values Method $\psi$
QP-GC Gurobi BM
$|\mathbb{J}^{\psi}_{p}|$ 6 0 2
$\bar{\mathbb{J}}^{\psi}$ 0.83 0.34 0.77
To conclude, we find that our method demonstrates the ability to recover clusters not only in synthetic graphs but also in real-world graphs.
5 Conclusion
In this paper, we considered a class of optimization models for graph clustering problems in a unified manner. We reformulated these models as sparse-constrained optimization problems and recovered the solutions of the original problem from the solutions of the relaxed problem without sparse constraints. A quadratic penalty method was applied to the relaxed problem. Our method QP-GC was compared with Gurobi and the Boltzmann machine. Through extensive numerical experiments, we found that, for both synthetic and real-world graphs, QP-GC not only outperformed the other methods in speed but also provided more accurate clustering results.